OpenStudy (anonymous):

What is an equation of a line that passes through the point (-2,5) and is perpendicular to the line whose equation is y= 1/2x+5

OpenStudy (anonymous):

1) Note that the gradient of y = 1/2(x) + 5 is 1/2. It is the number before the x. 2) Two perpendicular lines have gradients which multiply to give -1, so let 1/2m = -1. Then m = -2 3) Substitute into the following equation: y - b = m(x - a), for a line with gradient m, passing through (a,b). y - (5) = -2(x - (-2)) y - 5 = -2x - 4 y = -2x + 1

OpenStudy (anonymous):

Thank you so much you're a life saver!!!!

OpenStudy (anonymous):

No problem :)